OpenStudy (anonymous):

In one titration 20.00mL of the unknown sulfuric acid was placed into 125 mL Erlenmeyer flask with a drop of phenolphthalein solution and the 0.50M KOH added. A pink color appeared when 28.7 mL of the KOH had been added. What is the concentration of the sulfuric acid?

OpenStudy (anonymous):

use M1V1=M2V2; plug in your value inside such as; (0.02)(M)=(0.05)(0.0287) then you can find the conc. of H2SO4. Note: remember to change mL into L

OpenStudy (rogue):

Remember to keep in mind that sulfuric acid is diprotic (it can lose 2H+). So 2 KOH's are required to neutralize 1 H2SO4. If you are taking college chem/AP chem, you might have to consider the equilibrium for the 2nd proton, but this problem probably isn't meant to be that complicated, so don't worry about it.

OpenStudy (rogue):

\[H_2 SO_4 + 2 KOH \rightarrow 2H_2 O + K_2 SO_4\]\[M_1 V_1 = M_2 V_2\]\[M_1 \times 0.145 L = 0.50 M \times 0.0287 L\]\[M_1 = 0.09896 mol/L\]That is the value of M1 assuming it only loses 1 H+, but since it is losing 2 protons, the concentration of H2SO4 is half of that. Give it a little thought and it'll make sense.\[M_{H_2 SO_4} = 0.0494827 M\]